[12pt]article  Species Tag:  31010  Name:  CH_{3}CH_{2}D g.s and ν_{1}8 Version:  1  1D-Ethane Date:  Dec. 2014 Contributor:  A.M. Daly, B. J. Drouin Lines Listed: 38764  Q(300.0)= 32720.6379 Freq. (GHz) < 2000  Q(225.0)= 18267.0606 Max. J:  50  Q(150.0)= 8698.1224 LOGSTR0=  -12.0  Q(75.00)= 2858.7608 LOGSTR1=  -12.0  Q(37.50)= 1008.6979 Isotope Corr.: 0  Q(18.75)= 359.2054 Egy. (cm^{−1}) > 0.0  Q(9.375)= 128.8711 μ_{a} =  0.015 (gs)  A=  69653.392 μ_{b} =  0.015  B=  18859.082 μ_{c} =  C=  18214.160 Initial work by E. Hirota, Y. Endo, S. Saito, J.L. Duncan, Microwave spectra of deuterated ethanes: J. Mol. Spectrosc. 89 (1981) 285-295 measured the rotational spectrum in the ground state of CH_{3}CH_{2}D up to 160 GHz including 12 pairs of a− and 15 pairs of b− dipole transitions that were split into two hindered rotation components (A and E). Subsequent measurements made at JPL by A.M.Daly, B.J. Drouin, P. Groner, S.Yu and J. C. Pearson (JMS 307, p. 27-32, 2015) report the pure rotational spectrum of the ground and first excited torsional state ν_{1}8 of CH_{3}CH_{2}D, with measurements made up to 1.6 THz for the ground state and 1.1 THz for the ν_{1}8 state. The energy differences between the A and E torsional substates, ΔE(E-A), of 74.167(18) and -3382.23(34) MHz for the ground and excited states, respectively. Using these energy differences and the overtone transitions Δν_{18} = 2 from Raman measurements in the literature J.M. Fernández-Sánchez, S. Montero, J. Chem. Phys. 94 (1991) 7788-7800, the coefficients V_{3} and V_{6} of the potential function of the internal rotation in CH_{3}CH_{2}D were determined as V_{3}= 1004.56(4) cm^{−1} and V_{6}= 7.09(12) cm^{−1}. The dipole moment components are estimated to be equal components of the deuteromethane dipole (J. K. G. Watson, M. Takami, and T. Oka, 1979, J. Chem. Phys., 70, 5376). The partition function at 300 K was calculated with a sum-over-states up to J of 80, and does include other vibrational states below 990 cm^{−1}. The summation is truncated at J = 50.